OpenStudy (anonymous):

Can anyone explain why \( 1+2+3+4+...= -\dfrac{1}{12}\)?

OpenStudy (anonymous):

No, because it doesn't?

OpenStudy (anonymous):

Do you know zeta function?

OpenStudy (anonymous):

\[\Large\zeta(n) = \sum_{k=1}^\infty \dfrac{1}{k^n}\]But according to Wolfram Alpha, \(\zeta(-1) = -\dfrac{1}{12}\). http://www.wolframalpha.com/input/?i=zeta%28-1%29

OpenStudy (loser66):

Siths is not wrong, your question is ambiguous.

OpenStudy (anonymous):

Sorry, I know of it, but I'm not familiar with it, unfortunately. The way I read your series was \[1+2+3+\cdots=\sum_{n=1}^\infty n=\infty\] Which arguably is identical to the series \[\sum_{k=1}^\infty \frac{1}{k^n},\text{ with }n=-1~...\]

OpenStudy (anonymous):

Well so Wolfram Alpha is wrong?

OpenStudy (anonymous):

Not necessarily. I may be wrong in saying the two series are identical. You'll have to forgive my ignorance regarding the zeta function.